OpenStudy (anonymous):

A positive integer is picked randomly from 61 to 70, inclusive. What is the probability that it is prime?

OpenStudy (anonymous):

pls explain the concept of prime no

OpenStudy (anonymous):

The numbers from 61 to 70 are: 61, 62, 63, 64, 65, 66, 67, 68, 69, 70 Prime numbers: 61, 63, 67, 69

OpenStudy (anonymous):

do you know what "prime number " means?

OpenStudy (anonymous):

i m confused

OpenStudy (hoblos):

the prime numbers are 61 & 67 only

OpenStudy (anonymous):

lol prime is not odd!

OpenStudy (anonymous):

prime means only divisors are 1 and itself, a whole number that you cannot factor

OpenStudy (hoblos):

so the probability is 2/10 = 1/5

OpenStudy (anonymous):

we can try the numbers one by one and check

OpenStudy (anonymous):

okay,pratu u were mistaken

OpenStudy (anonymous):

61, 62, 63, 64, 65, 66, 67, 68, 69, 70 61 is prime because it is not the product of any two numbers. think about the multiplication table. 61 does not appear as an entry there

OpenStudy (anonymous):

62 is even, so it is divisible by 2, not prime. and neither are 64,66,68,70 for the same reason

OpenStudy (anonymous):

63 is divisible by 3 because \[63=3\times 21\]

OpenStudy (anonymous):

why did u take such a big no?

OpenStudy (anonymous):

65 is obviously divisible by 5 67 is prime and finally 69 is divisible by 3

OpenStudy (anonymous):

I think a better question is how do you know if a number is a prime. We know that all even numbers are not prime. But which of the odd ones are prime. I guess the question is how would you calculate if a number is or is not prime? For example, 3335679 ?

OpenStudy (anonymous):

true

OpenStudy (anonymous):

Can this only be done by brute force. i.e. take each integer and start dividing by the smaller integers and then keep going until you reach an answer without a remainder? We need a mathematician.

OpenStudy (anonymous):

well sorry i m a lost in maths

OpenStudy (anonymous):

Yeah, I made a mistake and had some internet problem in the middle so couldn't reply. The others are right though.

OpenStudy (hoblos):

some notes that may help: if the number is even then it is divisible by 2 if the sum of digits of the number is divisible by 3 then the whole number is divisible by 3 if the number ends with 0 or 5 then it is divisible by 5 out of these cases you have to try dividing by PRIME numbers until you reach an answer without a remainder